(a) . Given that in an economy

C = 0.8 (1-t) Y

t = 0.25

I = 900 - 50i

𝐺̅ = 800

L = 0.25Y - 62.5i

(𝑀̅/𝑃̅) = 500

(a) Derive the equation for the IS curve.

(c) What are the equilibrium levels of income and

the interest rate?

(d) Monetary & fiscal policy multiplier. 6

b) State whether the following statements are

TRUE or FALSE. Give reason(s) in support

of your answer. 5

i. Higher the marginal propensity to consume,

higher is the size of multiplier

ii. If investment is very sensitive to interest rate,

then we have a flat IS curve

(c) Use the following information (in rupees):

Income (Y) = 1,00,000

Nominal Money Supply (M) = 80,000

Price Level (P) = 20

Calculate the money growth rate required to

finance the budget deficit of Rs.10,000 in an

economy.


1
Expert's answer
2021-07-29T12:34:02-0400

A. (a) Equation of IS curve.

"c=0.8(1-t)Y"

"t=0.25"

"I=900-50i"

"L=0.25Y-62.5i"


"y=c+I"

"y=0.8(1-t)y+(900-50i)"

"y(0.8t-0.8)=900-50i"

"y=\frac {900-50i}{0.8t-0.8}"

(b) Equilibrium level of income

"Y=C+I+G"

"Y=0.8(1-t)Y+900-50i+800"

"Y=0.8(1-t)Y+1700-50i"

(c)

Fiscal multiplier

Given by "\frac{\Delta Y}{\Delta G}"

"=\frac{0.8(1-t)Y+1700-50i}{800}"

Monetary multiplier - driven by the central bank which controls the money supply via the interest rates.

monetary multiplier"=\frac{1}{Reserve Ratio}"

B.

1.TRUE - The higher the marginal prospensity to consume, the higher the size of the multiplier. This is because changes in income levels lead to proportionately larger changes in the consumption of a particular good.

2.TRUE- If investment is very sensitive to interest rate, then we have a flat IS curve. This is because a reduction in interest rate causes a big increase in national income and product, hence IS curve is flat.

C.

"Y=100,000"

"Money Supply(M)=80,000"

"Price Level (P)=20"

Money Growth Rate ="\frac {P\times Y}{M}"

"=\frac {20\times100,000}{80,000}"

=25%.